# Reducing Stochastic Games to Semidefinite Programming

### v1

@@ Abstract
We present a polynomial-time reduction from max-average constraints to the feasibility problem for semidefinite programs. This shows that Condon’s simple stochastic games, stochastic mean payoff games, and in particular mean payoff games and parity games can all be reduced to semidefinite programming.
@@ 1 Introduction
There are two important clusters of computational problems that are not known to be in the complexity class P: the first is related to numeric computation, and contains the sums-of-square-roots problem [GGJ76, EHS24], the Euclidean shortest path problem, PosSLP [ABKPM09], and the feasibility problem for semidefinite programs [NN94, Ram97]. The second cluster is related to tropical geometry, and contains for instance the model checking problem for the propositional $\mu$-calculus [EJ91, EJS93], parity games [Mar75, Mos91], mean payoff games [EM79], and/or scheduling [MSS04], stochastic mean payoff games [AM09], and simple stochastic games [Con92]. So far, no polynomial-time reduction from a problem of one of the clusters to a problem from the other cluster was known. We show that all of the mentioned problems from the second cluster can be reduced in polynomial time to the feasibility problem for semidefinite programs from the first cluster.
@@
Semidefinite programming is a generalisation of linear programming with many important algorithmic applications in discrete and continuous optimisation, both from a theoretical and a practical perspective [GLS94, BTN01, Las09]. However, already the feasibility problem for semidefinite programs is not known to be in P. By Ramana’s duality [Ram97], the problem is either in the intersection of NP and coNP, or outside of the union of NP and coNP. The semidefinite feasibility problem falls into the existential theory of the reals, which is known to be in PSPACE [Can88]. However, it has been observed by Khachiyan that the smallest feasible solution to an SDP might be of doubly exponential size (see, e.g., [PK97]), which is an obstacle for the polynomial-time algorithmic methods known for linear programming. In fact, it is possible to reduce the PosSLP problem to semidefinite programming [TV08]. PosSLP has been introduced in [ABKPM09], motivated by the ‘generic task of numerical computation’. For example, the sums-of-square-roots problem, which is a numeric computational problem not even known to be in NP and the barrier for polynomial-time tractability for numerous problems in computational geometry, such as the Euclidean shortest path problem, has a polynomial-time reduction to PosSLP [ABKPM09].
@@
Mean payoff games are turn-based two-player games on graphs [EM79, ZP96]; it is a famous open problem in theoretical computer science whether there is a polynomial-time algorithm to decide for a given graph which of the players has a winning strategy. Finding the winning region in mean payoff games is polynomial-time equivalent to various other computational problems in theoretical computer science, for instance scheduling with and-or-constraints [MSS04], the max-atoms problem [BNRC08], as well as solvability of max-plus systems and tropical linear feasibility [ABG07, AGG12]. The latter can be seen as the tropical analog of testing feasibility of linear inequalities (which for classical geometry is known to be in P, e.g. via the ellipsoid method [Kha79, GLS94]).
@@
Furthermore, there is a simple reduction from parity games to mean payoff games by using the priorities of the parity game with a suitably large basis [Pur95]. Parity games are polynomial-time equivalent to the model-checking problem of the propositional $\mu$-calculus [EJS93], which has been called ‘the most important logics in model checking’ [BW18]. For parity games, a quasipolynomial algorithm has been found recently [CJK+22]. Subsequent work indicates that the various quasipolynomial methods that have been obtained lately for parity games [LPSW22, JMT22, LB20] do not extend to mean payoff games [CFGO22].
[…]
@@
We introduce a fragment of max-plus-average constraints, which we call max-average constraints. There is a polynomial-time reduction from stopping simple stochastic games to max-average constraints (Section 3). Max-average constraints have the advantage that they can be further reduced to the feasibility problem for semidefinite programs. This implies that all the mentioned computational problems from the second cluster can be reduced to semidefinite programming. See Figure 1 for an overview of the mentioned computational problems and their relationship.
The reduction uses a similar idea as the reduction of Schewe [Sch09] to reduce parity games and mean payoff games to linear programming, which is, however, not a proper polynomial-time reduction. The idea is to replace constraints of the form $x_{0}=\max(x_{1},\dots,x_{k})$ by $x_{0}=\log_{b}(b^{x_{1}}+\dots+b^{x_{k}})$ for a sufficiently large $b$. Following [AGS18], we extend this idea to max-average constraints and obtain a proper polynomial-time reduction to non-Archimedean SDPs. From there, we use bounds from quantifier elimination results to translate non-Archimedean SDPs to real SDPs. However, we still have the obstacle that these SDPs involve coefficients of doubly exponential size, which means that they have exponential representation size and do not lead to a polynomial-time reduction. We overcome this difficulty by using the duality of semidefinite programming to find small expressions that define coefficients of doubly exponential size, which combined with the above ideas leads to a polynomial-time reduction from max-average constraints to (real) SDPs.
Our results imply that if there is a polynomial-time algorithm for SDPs, then all the mentioned problems in the second cluster can be solved in polynomial time as well. Conversely, our reduction can be used to translate interesting families of instances for (stochastic) mean payoff games or parity games into SDP instances, which might yield interesting families of instances for (algorithmic approaches to) semidefinite programming. For linear programming, this idea turned out to be very fruitful: certain instances of parity games have been used to prove exponential lower bounds for several famous pivoting rules for the simplex method in linear programming [FHZ11, Fri11, FHZ14, AF17, DFH23]. However, these results were not based on a general polynomial-time reduction from parity games to linear programming (which is not known!). We also hope that algorithmic ideas for solving games might generalise to (certain classes of) SDPs.
@@ 1.1 Related work
There are multiple works that reduce games to continuous optimization problems. As mentioned above, one-player mean payoff games (Markov decision processes) can be reduced to linear programming [Put05], which implies that they can be solved in polynomial time. For the case of two-player deterministic mean payoff games, Schewe [Sch09] proposed a reduction to linear programs with coefficients of exponential representation size. Such programs can also be interpreted as linear programs over non-Archimedean ordered fields [ABGJ15, ABGJ14, ABGJ21]. The idea of Schewe was generalized to stochastic games in [BEGM17], where the authors prove that stochastic mean payoff games can be encoded by convex programs with constraints of exponential encoding size. Such programs can be solved by the ellipsoid method in pseudo-polynomial time provided that the number of stochastic nodes is fixed. The non-Archimedean approach was generalized from deterministic games to stochastic games in the work [AGS18], which proves that stochastic mean payoff games can be reduced to semidefinite programs over non-Archimedean fields. The paper [AGS18] serves as a basis for our results.
These reductions express mean payoff games as convex optimization problems, but these problems are “non-standard” in the sense that they either use constraints of exponential encoding size or are expressed over non-Archimedean fields. There is yet another line of work deriving polynomial-time reductions from mean payoff games to classical optimization problems which are not convex. In particular, [Con93] expressed simple stochastic games as non-convex quadratic programs. This reduction is further studied in [KRSW22]. Another reduction to (generalized) complementarity problems with $P$-matrices was proposed in [GR05, SV06] and extended in subsequent works [JS08, FJS10, HIJ13]. Such linear complementarity problems form a subclass of quadratic programs and belong to the complexity class UniqueEOPL [FGMS20]. However, they are in general non-convex and can therefore not be expressed by semidefinite programs.
@@ 2 Preliminaries
The set of natural numbers $\{0,1,2,\dots\}$ is denoted by ${\mathbb{N}}$, and the set of real numbers by ${\mathbb{R}}$. We assume familiarity with first-order logic; see, e.g., [Hod97]. Let $\tau$ be a signature, i.e., a set of function and relation symbols. A first-order $\tau$-formula is a formula build from the Boolean connectives, atomic $\tau$-formulas and quantification in the usual way. If $A$ is a $\tau$-structure and $\phi$ is a $\tau$-formula with free variables $x_{1},\dots,x_{k}$, then $\phi$ defines over $A$ the relation
@@
A set if called semialgebraic if it has a first-order definition over the field $({\mathbb{R}};+,\cdot)$, allowing real parameters. A first-order formula $\phi(x_{1},\dots,x_{k})$ with free variables $x_{1},\dots,x_{k}$ is called primitive positive if it is of the form
@@ 2.1 Semidefinite Programming
For a symmetric matrix $S\in{\mathbb{R}}^{m\times m}$ we write $S\succeq 0$ if $S$ is positive semidefinite, i.e., if $x^{\top}Sx\geq 0$ for all $x\in{\mathbb{R}}^{m}$, and $S\succ 0$ if it is positive definite, i.e., if $x^{\top}Sx>0$ for all $x\in{\mathbb{R}}^{m}\setminus\{0\}$. A semidefinite program (SDP) is an optimisation problem of the following form:
@@
where $A_{1},\dots,A_{n},B\in{\mathbb{R}}^{m\times m}$ are symmetric and $c_{1},\dots,c_{n}\in{\mathbb{R}}$. If the data $A_{1},\dots,A_{n},B,c_{1},\dots,c_{n}$ is rational, then the problem of deciding whether the supremum exists, and if so, to compute the supremum, is a computational problem known as the semidefinite programming problem. The representation size of the SDP then consists of the sum of the bit lengths of all the rational numbers in the input.
The expression $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0$ in Section 2.1 is called a linear matrix inequality (LMI), and defines the feasible region of the SDP, which is called a spectrahedron. The representation size of LMIs is defined analogously as for SDPs. We say that an SDP is strictly feasible if $\sum_{i=1}^{n}x_{i}A_{i}-B\succ 0$ has a solution $x\in{\mathbb{R}}^{n}$.
@@
$x_{1}A_{1}+\dots+x_{n}A_{n}-B\succeq 0\enspace.$
Note that spectrahedra are convex and semialgebraic. It has been conjectured that every convex semialgebraic set is the projection of a spectrahedron (also called a spectrahedral shadow), but this turned out to be false [Sch18] (see [BKT24] for an alternative proof of an even stronger inexpressibility result). It is easy to see that the set of spectrahedral shadows is closed under primitive positive definability, and that for every primitive positive formula $\phi$ we may find an LMI for the spectrahedral shadow defined by $\phi$ whose representation size is bounded by the sum of the representation sizes of conjuncts of $\phi$. We also note that the definition of a semidefinite program makes sense over any real closed field, not only over the real numbers. In particular, it is meaningful to consider semidefinite programs over non-Archimedean real closed fields, such as the field of Puiseux series which we discuss in Section 4.
The (Frobenius) inner product of two matrices $S,T\in{\mathbb{R}}^{m\times m}$ is defined as $\langle S,T\rangle:=\text{trace}(ST)=\sum_{i,j}S_{ij}T_{ij}$. The dual program to the SDP from Section 2.1 (to which we refer as the primal) is given by
@@
It is well-known that the value of the primal SDP is bounded from below by the value of the dual; however, the lower bound might not be tight, i.e., there might be a non-zero duality gap. Moreover, both primal and dual might not have optimal solutions, i.e., the supremum may not be a maximum and the infimum might not be a minimum. See [Ram97] for a more complicated, but exact dual. However, for us the following sufficient condition for a vanishing duality gap will be enough.
Theorem 2.2 (see Theorem 2.4.1 in [BTN01]).
@@ 2.2 Stochastic Games
A stochastic mean payoff game is a two-player game played on a finite edge-weighted directed graph $\vec{\mathcal{G}}=(V,E,r)$, i.e., $V$ is a finite set of nodes, $E\subseteq V^{2}$ is a set of edges, and $r\colon E\to{\mathbb{Q}}$ is a weight function. The nodes of the graph are partitioned into three subsets, $V=V_{\mathrm{Min}}\sqcup V_{\mathrm{Max}}\sqcup V_{\mathrm{Rand}}$. The nodes in $V_{\mathrm{Min}}$ are controlled by player Min, the nodes in $V_{\mathrm{Max}}$ are controlled by player Max, and the nodes in $V_{\mathrm{Rand}}$ are controlled by nature. The players play the game by moving a pawn on the graph. When the pawn is on a node $v\in V_{\mathrm{Min}}$, player Min chooses an outgoing edge $(v,u)$ and moves the pawn to $u$. Analogously, player Max chooses the next move when the pawn is on a node $v\in V_{\mathrm{Max}}$. When the pawn is on a node $v\in V_{\mathrm{Rand}}$, the next move is chosen by nature randomly, according to some probability distribution on the outgoing edges of $v$. This probability distribution is known to both players and does not change during the game. To ensure that players can always make a move, we suppose that all nodes of the graph have at least one outgoing edge. The weights of the edges of the graph represent payoffs that player Max receives from player Min after each move of the pawn. In order to state the objectives of both players, we first specify the strategies that they are allowed to use. A pure and positional strategy (also called a policy) for player Min is a function $\sigma\colon V_{\mathrm{Min}}\to V$ that to each node controlled by Min associates a neighbor, $(v,\sigma(v))\in E$. Player Min plays according to $\sigma$ if they always move the pawn to $\sigma(v)$ when the pawn lands on $v$. We analogously define a policy $\tau\colon V_{\mathrm{Max}}\to V$ of player Max. Note that if we fix a pair of policies and an initial position of the pawn, then the resulting movement of the pawn is a Markov chain on the graph $(V,E)$. Given an initial position $u\in V$ of the pawn and a pair of policies $(\sigma,\tau)$, the mean payoff reward of player Max is the quantity
@@
where $u=u_{0},u_{1},\dots,u_{N}$ is the path of the pawn in the first $N$ turns of the game and the expectation is taken with respect to the Markov chain obtained by fixing $(\sigma,\tau)$. The ergodic theorem of Markov chains shows that $g(\sigma,\tau)_{u}$ is well-defined. The objective of player Max is to maximize the mean payoff reward while the objective of player Min is to minimize this reward. It is known that both players in a stochastic mean payoff game have optimal policies [LL69]. In other words, there exist policies $(\sigma^{*},\tau^{*})$ and a vector $\chi\in\mathbb{R}^{V}$ such that for every $u$ and every $(\sigma,\tau)$ we have the inequality
@@
A simple stochastic game is a special case of a stochastic mean payoff game. In this game, every node in $V_{\mathrm{Rand}}$ has out-degree exactly two and the probability distribution at every such node is $(1/2,1/2)$. Furthermore, the graph is equipped with two special terminal states called $\mathrm{Win}$ and $\mathrm{Lose}$. These states are absorbing, i.e., $\mathrm{Win}$ has only one outgoing edge that goes back to $\mathrm{Win}$ and $\mathrm{Lose}$ has only one outgoing edge that goes back to $\mathrm{Lose}$. The weight of the edge going from $\mathrm{Win}$ to $\mathrm{Win}$ is equal to $1$, while the weight of all of the other edges of the graph is equal to 00. In this case, the mean payoff reward of player Max is simply the probability that the pawn reaches the winning state, $g(\sigma,\tau)_{u}=\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win}))$. A simple stochastic game is called stopping if for every initial position $u\in V$ and every pair of policies $(\sigma,\tau)$, the pawn reaches a terminal state with probability one, $\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win},\mathrm{Lose}))=1$. In other words, the game is stopping if the Markov chain obtained by fixing any pair of policies has only two recurrent classes, namely $\{\mathrm{Win}\}$ and $\{\mathrm{Lose}\}$.
The computational problem associated with stochastic mean payoff games is the following: given a stochastic mean payoff game, find its value and a pair of optimal policies. In [AM09], Andersson and Miltersen proved that this problem is polynomial-time (Turing) reducible to the problem of finding the value of a stopping simple stochastic game. By a binary search argument, the problem of computing the value of a stopping simple stochastic game can be further reduced to the decision problem, in which we want to decide whether the value of a given state of such a game is at least $1/2$. The details of this reduction can be found, e.g., in [AGS18] or [Sko18, Section 7.1].
Theorem 2.3 ([AGS18, Lemma 48]).

### v2

# Reducing Stochastic Games to Semidefinite Program Feasibility

@@ Abstract
We present a polynomial-time reduction from max-plus-average constraints to the feasibility problem for semidefinite programs. This shows that Condon’s simple stochastic games, stochastic mean payoff games, and in particular mean payoff games and parity games can all be reduced to semidefinite programming.
keywords:
Mean-payoff games, stochastic games, semidefinite programming, max-plus-average constraints, max-atom problem
@@ 1 Introduction
There are two important clusters of computational problems that are not known to be in the complexity class P: the first is related to numeric computation, and contains the sums-of-square-roots problem [35, 27], the Euclidean shortest path problem, PosSLP [7], and the feasibility problem for semidefinite programs [51, 55]. The second cluster is related to tropical geometry, and contains for instance the model checking problem for the propositional $\mu$-calculus [29, 28], parity games [48, 50], mean payoff games [26], and/or scheduling [49], stochastic mean payoff games [8], simple stochastic games [23], tropically convex and max-plus-average constraints [17]. So far, no polynomial-time reduction from a problem of one of the clusters to a problem from the other cluster was known. We show that all of the mentioned problems from the second cluster can be reduced in polynomial time to the feasibility problem for semidefinite programs from the first cluster.
@@
Semidefinite programming is a generalisation of linear programming with many important algorithmic applications in discrete and continuous optimisation, both from a theoretical and a practical perspective [37, 12, 44]. However, already the feasibility problem for semidefinite programs is not known to be in P. By Ramana’s duality [55], the problem is either in the intersection of NP and coNP, or outside of the union of NP and coNP. The semidefinite feasibility problem falls into the existential theory of the reals, which is known to be in PSPACE [21]. However, it has been observed by Khachiyan that the smallest feasible solution to an SDP might be of doubly exponential size (see, e.g., [52]), which is an obstacle for the polynomial-time algorithmic methods known for linear programming. In fact, it is possible to reduce the PosSLP problem (testing positivity of a value computed by a Straight Line Program) to semidefinite programming [60]. PosSLP has been introduced in [7], motivated by the ‘generic task of numerical computation’. For example, the sums-of-square-roots problem, which is a numeric computational problem not even known to be in NP and the barrier for polynomial-time tractability for numerous problems in computational geometry, such as the Euclidean shortest path problem, has a polynomial-time reduction to PosSLP [7].
@@
Mean payoff games are turn-based two-player games on graphs [26, 61]; it is a famous open problem in theoretical computer science whether there is a polynomial-time algorithm to decide for a given graph which of the players has a winning strategy. Finding the winning region in mean payoff games is polynomial-time equivalent to various other computational problems in theoretical computer science, for instance scheduling with and-or-constraints [49], the max-atoms problem [14], as well as solvability of max-plus systems and tropical linear feasibility [2, 1]. The latter can be seen as the tropical analog of testing feasibility of linear inequalities (which for classical geometry is known to be in P, e.g., via the ellipsoid method [42, 37]).
@@
Furthermore, there is a simple reduction from parity games to mean payoff games by using the priorities of the parity game with a suitably large basis [53]. Parity games are polynomial-time equivalent to the model-checking problem of the propositional $\mu$-calculus [28], which has been called ‘the most important logics in model checking’ [19]. For parity games, a quasipolynomial algorithm has been found recently [20]. Subsequent work indicates that the various quasipolynomial methods that have been obtained lately for parity games [46, 40, 45] do not extend to mean payoff games [22].
Mean payoff games can be generalised to stochastic mean payoff games (sometimes called $2\frac{1}{2}$-player games), where the graph might contain stochastic nodes. If the strategy of one of the two players in such a game is fixed, the game turns into a Markov decision process, for which polynomial-time algorithms based on linear programming are known [54]. Stochastic mean payoff games are equivalent under polynomial-time Turing reductions to Condon’s simple stochastic games [8] (we may even assume that the simple stochastic games are stopping, see the discussion in Section˜2.2), which are known to be in the intersection of NP and coNP [23]. There are many other variants of games that all have the same complexity, see [8].
One approach to analyze mean payoff games consists in reducing these games to constraint satisfaction problems (CSPs). More precisely, it is known that mean payoff games are equivalent to the max-atom problem [9, 49], while the more expressive stochastic mean payoff games can be reduced to max-plus-average constraints, which are still in NP $\cap$ coNP [17]. In the conference version of this article, we introduced a fragment of max-plus-average constraints, called max-average constraints. There is a polynomial-time reduction from stopping simple stochastic games to max-average constraints (Section˜3). Max-average constraints have the advantage that they can be further reduced to the feasibility problem for semidefinite programs, as we will show here. In fact, with some extra effort we can also show this for the entire class of max-plus-average constraints. This implies that all the mentioned computational problems from the second cluster can be reduced to semidefinite programming. See Figure˜1 for an overview of the mentioned computational problems and their relationship.
@@
Our reduction consists of two steps. The first step uses a similar idea as the reduction of Schewe [57] to reduce parity games and mean payoff games to linear programming, which is, however, not a proper polynomial-time reduction. More precisely, the idea of this first step is to replace constraints of the form $x_{0}=\max(x_{1},\dots,x_{k})$, which occur in the max-atom problem, by $x_{0}=\log_{b}(b^{x_{1}}+\dots+b^{x_{k}})$ for a sufficiently large $b$. This reduction is not a polynomial-time reduction since the numbers $b^{x_{i}}$ may have exponential bit-sizes for any $x$ that solves the original problem. Following [6], we extend this idea to max-plus-average constraints and obtain a proper polynomial-time reduction to non-Archimedean SDPs, which are SDPs defined over a field of formal power series. From there, we use bounds from quantifier elimination results to translate non-Archimedean SDPs to real SDPs. However, we still have the obstacle that these SDPs involve coefficients of doubly exponential size, which means that they have exponential representation size and do not lead to a polynomial-time reduction. In the second step of our reduction, we overcome this difficulty by using the duality of semidefinite programming to find small expressions that define coefficients of doubly exponential size, which combined with the above ideas leads to a polynomial-time reduction from max-plus-average constraints to (real) SDPs.
Our results imply that if there is a polynomial-time algorithm for SDPs, then all the mentioned problems in the second cluster can be solved in polynomial time as well. Conversely, our reduction can be used to translate interesting families of instances for (stochastic) mean payoff games or parity games into SDP instances, which might yield interesting families of instances for (algorithmic approaches to) semidefinite programming. For linear programming, this idea turned out to be very fruitful: certain instances of parity games have been used to prove exponential lower bounds for several famous pivoting rules for the simplex method in linear programming [33, 32, 34, 10, 25]. However, these results were not based on a general polynomial-time reduction from parity games to linear programming (which is not known!). We also hope that algorithmic ideas for solving games might generalise to (certain classes of) SDPs.
@@ 1.1 Related work
There are multiple works that reduce games to continuous optimization problems. As mentioned above, one-player mean payoff games (Markov decision processes) can be reduced to linear programming [54], which implies that they can be solved in polynomial time (but it is open whether there is a strongly polynomial algorithm). For the case of two-player deterministic mean payoff games, Schewe [57] proposed a reduction to linear programs with coefficients of exponential representation size. Such programs can also be interpreted as linear programs over non-Archimedean ordered fields [4, 3, 5]. The idea of Schewe was generalized to stochastic games in [18], where the authors prove that stochastic mean payoff games can be encoded by convex programs with constraints of exponential encoding size. Such programs can be solved by the ellipsoid method in pseudo-polynomial time provided that the number of stochastic nodes is fixed. The non-Archimedean approach was generalized from deterministic games to stochastic games in the work [6], which proves that stochastic mean payoff games can be reduced to semidefinite programs over non-Archimedean fields. The paper [6] serves as a basis for our results.
These reductions express mean payoff games as convex optimization problems, but the resulting problems are “non-standard” in the sense that they either use constraints of exponential encoding size or are expressed over non-Archimedean fields. There is yet another line of work deriving polynomial-time reductions from mean payoff games to classical optimization problems which are not convex. In particular, [24] expressed simple stochastic games as non-convex quadratic programs. This reduction is further studied in [43]. Another reduction to (generalized) complementarity problems with $P$-matrices was proposed in [36, 59] and extended in subsequent works [41, 31, 38]. Such linear complementarity problems form a subclass of quadratic programs and belong to the complexity class UniqueEOPL [30]. However, they are in general non-convex and can therefore not be expressed by semidefinite programs.
@@ 2 Preliminaries
The set of natural numbers $\{0,1,2,\dots\}$ is denoted by $\mathbb{N}$}, and the set of real numbers by $\mathbb{R}$}. We assume familiarity with first-order logic; see, e.g., [39]. Let $\tau$ be a signature, i.e., a set of function and relation symbols. A first-order $\tau$-formula is a formula built from the Boolean connectives, atomic $\tau$-formulas and quantification in the usual way. If $A$ is a $\tau$-structure and $\phi$ is a $\tau$-formula with free variables $x_{1},\dots,x_{k}$, then $\phi$ defines over $A$ the relation
@@
A set is called semialgebraic if it has a first-order definition over the field $({\mathbb{R}};+,\cdot)$, allowing real parameters (we do not need the order in the signature since it is definable from addition and multiplication). A first-order formula $\phi(x_{1},\dots,x_{k})$ with free variables $x_{1},\dots,x_{k}$ is called primitive positive if it is of the form
@@ 2.1 Semidefinite Programming
For a matrix $S\in{\mathbb{R}}^{m\times m}$ we write
  • $S\succeq 0$ if $S$ is positive semidefinite, i.e., $S$ is symmetric and $x^{\top}Sx\geq 0$ for all $x\in{\mathbb{R}}^{m}$, and
  • $S\succ 0$ if it is positive definite, i.e., $S$ is symmetric and $x^{\top}Sx>0$ for all $x\in{\mathbb{R}}^{m}\setminus\{0\}$.
A semidefinite program (SDP) is an optimisation problem of the following form:
@@
where $A_{1},\dots,A_{n},B\in{\mathbb{R}}^{m\times m}$ are symmetric and $c_{1},\dots,c_{n}\in{\mathbb{R}}$. If the data $A_{1},\dots,A_{n},B$, $c_{1},\dots,c_{n}$ is rational, then the problem of deciding whether the infimum exists, and if so, to compute the infimum, is a computational problem known as the semidefinite programming problem. The representation size of the SDP then consists of the sum of the bit lengths of all the rational numbers in the input.
The expression $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0$ in Section˜2.1 is called a linear matrix inequality (LMI), and defines the feasible region of the SDP, which is called a spectrahedron. The representation size of LMIs is defined analogously as for SDPs. We say that an SDP is strictly feasible if $\sum_{i=1}^{n}x_{i}A_{i}-B\succ 0$ has a solution $x\in{\mathbb{R}}^{n}$.
@@
$x_{1}A_{1}+\dots+x_{n}A_{n}-B\succeq 0\kern 5.0pt.$
Note that spectrahedra are convex and semialgebraic. It has been conjectured that every convex semialgebraic set is the projection of a spectrahedron (also called a spectrahedral shadow), but this turned out to be false [56] (see [15] for an alternative proof of an even stronger inexpressibility result). It is easy to see that the set of spectrahedral shadows is closed under primitive positive definability, and that for every primitive positive formula $\phi$ we may find an LMI for the spectrahedral shadow defined by $\phi$ whose representation size is bounded by the sum of the representation sizes of conjuncts of $\phi$. We also note that the definition of a semidefinite program makes sense over any real closed field, not only over the real numbers. In particular, it is meaningful to consider semidefinite programs over non-Archimedean real closed fields, such as the field of Puiseux series which we discuss in Section˜4.
The (Frobenius) inner product of two matrices $S,T\in{\mathbb{R}}^{m\times m}$ is defined as $\langle S,T\rangle:=\text{trace}(ST)=\sum_{i,j}S_{ij}T_{ij}$. The dual program to the SDP from Section˜2.1 (to which we refer as the primal) is given by
@@
It is well-known that the value of the primal SDP is bounded from below by the value of the dual; however, the lower bound might not be tight, i.e., there might be a non-zero duality gap. Moreover, both primal and dual might not have optimal solutions, i.e., the supremum may not be a maximum and the infimum might not be a minimum. See [55] for a more complicated, but exact dual. However, for us the following sufficient condition for a vanishing duality gap will be enough.
Theorem 2.2 (see Theorem 2.4.1 in [12] and the explanations on page 142 and 143).
@@ 2.2 Stochastic Games
A stochastic mean payoff game is a two-player game played on a finite edge-weighted directed graph $\vec{\mathcal{G}}=(V,E,r,p)$, i.e., $V$ is a finite set of nodes, $E\subseteq V^{2}$ is a set of edges, $r\colon E\to{\mathbb{Q}}$ is a weight function, and $p\colon E\to{\mathbb{Q}}\cap[0,1]$ specifies edge probabilities. The nodes of the graph are partitioned into three subsets, $V=V_{\mathrm{Min}}\sqcup V_{\mathrm{Max}}\sqcup V_{\mathrm{Rand}}$. The nodes in $V_{\mathrm{Min}}$ are controlled by player Min, the nodes in $V_{\mathrm{Max}}$ are controlled by player Max, and the nodes in $V_{\mathrm{Rand}}$ are controlled by nature; the probabilities of the outgoing edges of each vertex in $V_{\mathrm{Rand}}$ sum up to one. The players play the game by moving a pawn on the graph. When the pawn is on a node $v\in V_{\mathrm{Min}}$, player Min chooses an outgoing edge $(v,u)$ and moves the pawn to $u$. Analogously, player Max chooses the next move when the pawn is on a node $v\in V_{\mathrm{Max}}$. When the pawn is on a node $v\in V_{\mathrm{Rand}}$, the next move is chosen by nature randomly according to the probability distribution given by $p$. To ensure that players can always make a move, we suppose that all nodes of the graph have at least one outgoing edge. The weights of the edges of the graph represent payoffs that player Max receives from player Min after each move of the pawn. In order to state the objectives of both players, we first specify the strategies that they are allowed to use. A pure and positional strategy (also called a policy) for player Min is a function $\sigma\colon V_{\mathrm{Min}}\to V$ that to each node controlled by Min associates a neighbor, $(v,\sigma(v))\in E$. Player Min plays according to $\sigma$ if they always move the pawn to $\sigma(v)$ when the pawn lands on $v$. We analogously define a policy $\tau\colon V_{\mathrm{Max}}\to V$ of player Max. Note that if we fix a pair of policies and an initial position of the pawn, then the resulting movement of the pawn is a Markov chain on the graph $(V,E)$. Given an initial position $u\in V$ of the pawn and a pair of policies $(\sigma,\tau)$, the mean payoff reward of player Max is the quantity
@@
where $u=u_{0},u_{1},\dots,u_{N}$ is the path of the pawn in the first $N$ turns of the game and the expectation is taken with respect to the Markov chain obtained by fixing $(\sigma,\tau)$. The ergodic theorem of Markov chains shows that $g(\sigma,\tau)_{u}$ is well-defined. The objective of player Max is to maximize the mean payoff reward while the objective of player Min is to minimize this reward. It is known that both players in a stochastic mean payoff game have optimal policies [47]. In other words, there exist policies $(\sigma^{*},\tau^{*})$ and a vector $\chi\in\mathbb{R}^{V}$ such that for every $u$ and every $(\sigma,\tau)$ we have the inequality
@@
A simple stochastic game is a special case of a stochastic mean payoff game. In this game, every node in $V_{\mathrm{Rand}}$ has out-degree exactly two and the probability distribution at every such node is $(1/2,1/2)$. Furthermore, the graph is equipped with two special terminal states called $\mathrm{Win}$ and $\mathrm{Lose}$. These states are absorbing, i.e., $\mathrm{Win}$ has only one outgoing edge that goes back to $\mathrm{Win}$ and $\mathrm{Lose}$ has only one outgoing edge that goes back to $\mathrm{Lose}$. The weight of the edge going from $\mathrm{Win}$ to $\mathrm{Win}$ is equal to $1$, while the weight of all of the other edges of the graph is equal to 0. In this case, the mean payoff reward of player Max is simply the probability that the pawn reaches the winning state, $g(\sigma,\tau)_{u}=\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win}))$.
A simple stochastic game is called stopping if for every initial position $u\in V$ and every pair of policies $(\sigma,\tau)$, the pawn reaches a terminal state with probability one, $\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win},\mathrm{Lose}))=1$. In other words, the game is stopping if the Markov chain obtained by fixing any pair of policies has only two recurrent classes, namely $\{\mathrm{Win}\}$ and $\{\mathrm{Lose}\}$.
The computational problem associated with stochastic mean payoff games is the following: given a stochastic mean payoff game, find its value and a pair of optimal policies. In [8], Andersson and Miltersen proved that this problem is polynomial-time (Turing) reducible to the problem of finding the value of a stopping simple stochastic game. By a binary search argument, the problem of computing the value of a stopping simple stochastic game can be further reduced to the decision problem, in which we want to decide whether the value of a given state of such a game is at least $1/2$. The details of this reduction can be found, e.g., in [6] or [58, Section 7.1].
Theorem 2.3 ([6, Lemma 48]).